# VP-MEL: Visual Prompts Guided Multimodal Entity Linking

### v2

@@ Abstract
Multimodal Entity Linking (MEL) is extensively utilized in the domains of information retrieval. However, existing MEL methods typically utilize mention words as mentions for retrieval. This results in a significant dependence of MEL on mention words, thereby constraining its capacity to effectively leverage information from both images and text. In situations where mention words are absent, MEL methods struggle to leverage image-text pairs for entity linking. To solve these issues, we introduce a Visual Prompts guided Multimodal Entity Linking (VP-MEL) task. VP-MEL directly marks specific regions within the image. These markers are referred to as visual prompts in VP-MEL. Without mention words, VP-MEL aims to utilize marked image-text pairs to align visual prompts with specific entities in the knowledge bases. A new dataset for the VP-MEL task, VPWiki, is proposed in this paper. Moreover, we propose a framework named FBMEL, which enhances the significance of visual prompts and fully leverages the information in image-text pairs. Experimental results on the VPWiki dataset demonstrate that FBMEL outperforms baseline methods across multiple benchmarks for the VP-MEL task.
@@
Hongze Mi1, Jinyuan Li2, Xuying Zhang1, Haoran Cheng1, Jiahao Wang1, Di Sun3, Gang Pan1,2,thanks:   corresponding authors. 1College of Intelligence and Computing, Tianjin University 2School of New Media and Communication, Tianjin University 3Tianjin University of Science and Technology {mhzqs, jinyuanli, wjhwtt, pangang}@tju.edu.cn dsun@tust.edu.cn
@@ 1 Introduction
Figure 1: Comparison between MEL and VP-MEL task. In the absence of mention words, MEL faces challenges in performing entity linking for the image content. VP-MEL overcomes this limitation by utilizing visual prompts to link specific regions in the images to the correct entities in knowledge base.
@@
Multimodal entity linking (MEL) (Moon et al., 2018) task aims to utilize both image and text information to disambiguate the mention and link it to the correct entity in the knowledge base(KB). Most MEL works (Gan et al., 2021; Wang et al., 2022a; Dongjie and Huang, 2022; Luo et al., 2023; Zhang et al., 2023a; Xing et al., 2023; Shi et al., 2024) mainly focus on improving the interaction of multimodal information and achieve promising performance. Existing methods represent mentions in the form of mention words and typically assume that each mention is associated with a high-quality image, which results in two limitations for MEL.
@@
Text information dependency: MEL relies on selecting mention words for entity linking, which often exhibit significant overlap with entity names in practical applications. This inherent similarity enables mention words to serve as strong cues for identifying specific entities within the knowledge base (KB). Although some ambiguity remains, this overlap substantially simplifies the entity linking task. When mention words are unavailable, MEL struggles to operate effectively and may produce erroneous outputs. As shown in Figure 1, the expected mention is Ted Kennedy on the right, but MEL generates incorrect responses, like Lolo Soetoro and Hartford. Many results, such as Lolo Soetoro, are entirely disconnected from both the image and text. Excessive reliance on mention words confines MEL to a text-focused task. It is necessary to introduce a new MEL task to overcome such constraints, enabling MEL to broaden its application scenarios and adapt to the multimodal era.
@@
Image modality impurity: Compared to text modality, image modality contains more noise. As shown in Figure 1, only the person on the right is relevant, while the one on the left is noise. Incorrect use of image information can significantly impact the results. Most existing coarse-grained methods (Yang et al., 2023; Luo et al., 2023; Wang et al., 2022a) encode the entire image directly, making it difficult to eliminate noise interference. Song et al. (2024) attempts to enhance visual features by utilizing fine-grained image attributes, but it heavily depends on high-quality images and may increase model size and training costs. To utilize images more effectively, it is necessary for MEL to find a way to overcome noise interference in images and focus on the regions related to the mention.
@@
Existing methods demonstrate limited effectiveness in utilizing image information and are overly dependent on mention words as textual cues. In many application scenarios, the inability of users to provide mention words significantly increases the difficulty of MEL. So we ask: Is it possible to introduce a MEL task that does not depend on mention words, thereby avoiding excessive reliance on text and enhancing the performance of image-based reasoning?
@@
In this paper, we introduce a new Visual Prompts guided Multimodal Entity Linking (VP-MEL) task. As shown in Figure 1, VP-MEL still focuses on addressing the entity linking task for image-text pairs. In the absence of mention words, VP-MEL uses a straightforward method to directly annotate mentions onto images using visual prompts, without any additional complex steps. Free from the constraints of mention words, VP-MEL supports a broader range of applications. It enables effective entity linking even in situations where users are unfamiliar with the textual language or place greater emphasis on image-based information. Moreover, a new VPWiki dataset is constructed based on the existing MEL public datasets, where visual prompts are annotated for each mention within the associated images.
@@
    We construct a high-quality annotated dataset, VPWiki, to enhance the benchmark evaluation of the task. A pipeline is proposed to efficiently automate data annotation for the VP-MEL task.
@@ 2.1 Multimodal Entity Linking
Moon et al. (2018) incorporate images into EL task to address the ambiguity and incompleteness of mentions in social media. Adjali et al. (2020) construct a MEL dataset of social media posts from Twitter and propose a model for jointly learning a representation of both mentions and entities from their textual and visual contexts. Gan et al. (2021) construct a dataset that contains long movie reviews with various related entities and images. Wang et al. (2022c) present a high-quality human-annotated MEL dataset with diversified contextual topics and entity types from Wikinews. In recent years, a multitude of outstanding works in the MEL field (Wang et al., 2022a; Yang et al., 2023; Luo et al., 2023; Shi et al., 2024; Song et al., 2024) have emerged, focusing on the utilization and optimization of multimodal information based on these public datasets. Although multimodal information can enhance entity linking performance, in these methods, text consistently dominates over images.
@@ 2.2 Vision Prompt
Region-specific comprehension in complex visual scenes has become a key research topic in the field of Multimodal Computer Vision. Existing methods utilize textual coordinate representations (Zhu et al., 2024; Zhao et al., 2023), learned positional embeddings (Peng et al., 2024; Zhang et al., 2023b; Zhou et al., 2023), or Region of Interest (ROI) features (Zhang et al., 2023b) to anchor language to specific image regions. Cai et al. (2024) propose a user-friendly visual prompting solution that directly overlays visual prompts onto the image canvas in a natural and intuitive way. Inspired by this solution, we introduce visual prompts into MEL to reduce the interference of noise in images and enhance understanding of specific regions.
@@ 3 Dataset
To address the VP-MEL task, we construct a VPWiki dataset as follows:
@@
Automated Annotation Pipeline.
To improve data annotation efficiency, we develop a pipeline that annotates visual prompts in images based on mention words, using the Visual Entailment Module and Visual Grounding Module, inspired by Li et al. (2024). This pipeline can automatically annotate a large amount of high-quality data, significantly reducing the manual annotation workload for subsequent tasks. The details of modules in the pipeline are provided in the Appendix A.6.
Manual Annotation Design.
Building on the WikiDiverse and WikiMEL datasets, annotators are provided with an image-text pair and the corresponding mention words. Annotators are required to: 1) annotate relevant visual prompts in the image based on the mention words; 2) remove samples where the image and mention words are unrelated; 3) re-annotate samples with inaccurate pipeline annotations; 4) annotate the entity type for each example (i.e., Person, Organization, Location, Country, Event, Works, Misc).
Figure 2: An example from VP-MEL and MEL. The red marker in the image represents the visual prompt annotated for the VP-MEL task. The red text indicates the annotated mention words.
Manual Annotation Procedure.
The annotators include 10 annotators and 2 experienced experts. All annotators have linguistic knowledge and are instructed with detailed annotation principles. Fleiss Kappa score (Fleiss, 1971) of annotators is 0.83, indicating strong agreement among them. We use the Intersection over Union (IoU) metric to assess annotation quality and discard samples with an IoU score below 0.5.
@@
Figure 3: More statistics of VPWiki. (a) Entity type distribution. (b) Distribution of the number of candidate entities per mention.
@@
Figure 2 presents an example from the VP-MEL dataset and compares it with MEL data. In total, 12,720 samples are obtained. After random sampling, the VPWiki dataset is split into training set, validation set, and test set with the ratio of 8:1:1. The detailed data statistics of VPWiki is illustrated in Table 1. The distribution of entity types and the number of candidate entities per mention in the VPWiki dataset are shown in Figure 3. Figure 3(a) uses abbreviations of each entity type for representation. In Figure 3(b), as the number of candidate entities per mention increases, the more difficult the task is.
@@ 4 Methodology
In this section, we first formulate the task of VP-MEL, and then go through the details of the proposed Feature Balanced Multimodal Entity Linking (FBMEL) framework.
4.1 Task Formulation
Related mathematical notations are defined as follows. The multimodal knowledge base is constructed by a set of entities =$\mathcal{E}=\{E_{i}\}_{i=1}^{N}$, and each entity is denoted as $E_{i}=(e_{n_{i}},e_{v_{i}},e_{d_{i}},e_{a_{i}})$, where the elements of $E_{i}$ represent entity name, entity images, entity description, and entity attributes, respectively. The mention is denoted as $M_{j}=(m_{s_{j}},m_{v_{j}})$, where $m_{s_{j}}$, $m_{v_{j}}$ indicate the sentence and the corresponding image. The related entity of mention $M_{j}$ in the knowledge base is $E_{i}$.
The task of VP-MEL targets to retrieve the ground truth entity $E_{i}$ from the entity set $\mathcal{E}$ of knowledge base based on $M_{j}$.
Figure 4: The overall architecture of Feature Balanced Multimodal Entity Linking (FBMEL) framework. The image-text pairs of the Mention and Entity are used together as input. Specifically, Mention Text is the sentence corresponding to Mention Image, while Entity Text consists of Entity Name and Entity Attribute corresponding to the Entity Image in Knowledge Base.
@@ 4.2 Visual Encoder
We choose to use pre-trained CLIP visual encoder (Dosovitskiy et al., 2021) to extract image features. The image $m_{v_{j}}$ of $M_{j}$ is reshaped into 2D patches, where the number of patches $n=H\times W/P^{2}$, with $H\times W$ representing the size of the image and $P$ representing the size of each patch. After this, image patches are processed through visual encoder to extract features. The hidden states extracted from $m_{v_{j}}$ by the CLIP visual encoder are represented as $V_{M_{j}}^{l}=\left[v_{[CLS]}^{0};v_{M_{j}}^{1};v_{M_{j}}^{2};...;v_{M_{j}}^{n}\right]\in\mathbb{R}^{(n+1)\times d_{c}}$, where $d_{c}$ denotes the dimension of the hidden state and $l$ denotes the number of layers in the encoder.
@@ 4.3 Detective-VLM
Background: $\{Image\}$ Text: $\{Sentence\}$ Question: In the red box of the image, tell me briefly what is the $\{Entity\ \allowbreak Type\}$, and which $\{Entity\ \allowbreak Name\}$ in the text corresponds to the $\{Entity\ \allowbreak Type\}$? Answer: $\{Entity\ \allowbreak Type\}$$\{Entity\ \allowbreak Name\}$
We utilize VPWiki dataset to design the fine-tuning dataset, where $\{Image\}$ and $\{Sentence\}$ correspond to $m_{v_{j}}$ and $m_{s_{j}}$ in $M_{j}$, respectively. During the inference process, $\{Entity\ \allowbreak Name\}$ and $\{Entity\ \allowbreak Type\}$ need to be generated by VLM, where the $\{Entity\ \allowbreak Type\}$ is one of [$Person$, $Organization$, $Location$, $Country$, $Event$, $Works$, $Misc$]. Details of the dataset and model for Detective-VLM can be found in Appendix A.4.
@@ 5.1 Experimental Settings
All the training and testing are conducted on a device equipped with 4 Intel(R) Xeon(R) Platinum 8380 CPUs and 8 NVIDIA A800-SXM4-80GB GPUs. Detailed experimental settings are provided in Appendix A.1. In order to conduct a comprehensive assessment of the efficacy of our approach, we compare FBMEL with various competitive MEL baselines and VLM baselines. Detailed introduction of baselines is shown in the Appendix A.2.
Datasets.
The VP-MEL experiments are based on VPWiki dataset and MEL experiments are based on the WikiDiverse (Wang et al., 2022c) dataset. Additional experiments and detailed explanations are provided in the Appendix A.8.
@@ 5.2 Main Results
In Table 2(a), FBMEL significantly outperforms the other methods. First, among the VLM methods, LLaVA-1.5 is the most competitive, showing differences of 5.14%, 1.97%, and 10.58% from FBMEL across the three metrics. Nonetheless, given the substantial data and computational resources required for training LLaVA, our method still demonstrates a clear advantage. Second, the MEL methods exhibit a performance gap compared to our method. MEL methods face challenges in performing effective entity linking when mention words are absent.
@@
Table 2(b) presents the experimental comparison results between FBMEL and the MEL baselines. During the training process, the Detective-VLM is removed from FBMEL, and its output is replaced with mention words from WikiDiverse. The metrics of our method are second only to MIMIC and MELOV, with a difference of only 0.27%, 0.28%, and 0.07% compared to MIMIC. FBMEL increases the emphasis on image features while reducing the influence of text features. It is reasonable that FBMEL performs slightly worse than MEL methods in the MEL task. Furthermore, the absence of visual prompts in images somewhat limits the ability of FBMEL to fully utilize visual information.

### v3

@@ Abstract
Multimodal entity linking (MEL), a task aimed at linking mentions within multimodal contexts to their corresponding entities in a knowledge base (KB), has attracted much attention due to its wide applications in recent years. However, existing MEL methods often rely heavily on mention words as retrieval cues, which limits their ability to effectively utilize information from both images and text. This reliance poses significant challenges in scenarios where mention words are absent, as current MEL approaches struggle to leverage image-text pairs for accurate entity linking. To solve these issues, we introduce a Visual Prompts guided Multimodal Entity Linking (VP-MEL) task. Given a text-image pair, VP-MEL aims to link a marked region (i.e., visual prompt) in an image to its corresponding entities in the knowledge base. To facilitate this task, we present a new dataset, VPWiki, specifically designed for VP-MEL. Furthermore, we propose a framework named FBMEL, which enhances visual feature extraction using visual prompts and leverages the pretrained Detective-VLM model to capture latent information. Experimental results on the VPWiki dataset demonstrate that FBMEL outperforms baseline methods across multiple benchmarks for the VP-MEL task. Code and datasets will be released at https://anonymous.4open.science/r/VP-MEL6-F9B0.
@@
Hongze Mi1, Jinyuan Li2, Xuying Zhang1, Haoran Cheng1, Jiahao Wang1, Di Sun3, Gang Pan1,2,thanks:   corresponding author. 1College of Intelligence and Computing, Tianjin University 2School of New Media and Communication, Tianjin University 3College of Artificial Intelligence, Tianjin University of Science and Technology {mhzqs, jinyuanli, wjhwtt, pangang}@tju.edu.cn
@@ 1 Introduction
Figure 1: Comparison between MEL and VP-MEL tasks. MEL struggles with entity linking for image content when mention words are unavailable. In contrast, VP-MEL addresses this limitation by using visual prompts to link specific regions in the images to the correct entities in the knowledge base.
@@
Linking ambiguous mentions with multimodal contexts to the referent unambiguous entities in a knowledge base (KB), known as Multimodal Entity Linking (MEL) (Moon et al., 2018), is an essential task for many multimodal applications. Most MEL works (Gan et al., 2021; Wang et al., 2022a; Dongjie and Huang, 2022; Luo et al., 2023; Zhang et al., 2023a; Xing et al., 2023; Shi et al., 2024) mainly focus on improving the interaction of multimodal information and achieve promising performance. However, existing methods typically represent mentions in the form of mention words and assume that each mention is associated with a high-quality image, which results in two limitations for MEL.
@@
Text information dependency: MEL heavily relies on mention words for entity linking, which often exhibit substantial overlap with entity names in practical applications. This inherent similarity enables mention words to serve as strong cues for identifying specific entities within the knowledge base (KB). Although some ambiguity remains, this overlap substantially simplifies the entity linking process. However, MEL struggles to perform effectively and often produces incorrect outputs when mention words are unavailable. As shown in Figure 1, the expected mention is Ted Kennedy (on the right), but MEL generates incorrect responses, like Lolo Soetoro and Hartford. Many results, such as Lolo Soetoro, are entirely unrelated to both the image and text. This excessive reliance on mention words confines MEL to a text-focused task. It is necessary to introduce a new MEL task to overcome such constraints, allowing MEL to broaden its application scenarios and adapt to the multimodal era.
@@
Image modality impurity: Compared to the text modality, the image modality often contains more noise. As illustrated in Figure 1, only the person on the right is relevant, while the one on the left represents noise. Misinterpreting or misusing such image information can seriously affect the results. Most existing coarse-grained methods (Yang et al., 2023; Luo et al., 2023; Wang et al., 2022a) encode the entire image directly, making it difficult to eliminate noise interference. Song et al. (2024) attempts to enhance visual features by utilizing fine-grained image attributes, but they heavily depends on high-quality images and often increase model size and training costs. To better leverage image information, it is necessary for MEL to find a way to overcome noise interference in images and focus on the regions related to the mentions.
@@
Existing methods demonstrate limited effectiveness in utilizing image information and are overly dependent on mention words as textual cues. In many application scenarios, the challenge of MEL is considerably worsened by the inability of users to provide relevant mention words. So we ask: Is it possible to introduce a MEL task that does not depend on mention words, thereby avoiding excessive reliance on text and enhancing the performance of image-based reasoning?
@@
In this paper, we define a new Visual Prompts guided Multimodal Entity Linking (VP-MEL) task. As shown in Figure 1, VP-MEL is designed to address the entity linking task for image-text pairs. VP-MEL employs a straightforward method to annotate mentions directly onto images using visual prompts, eliminating the need for mention words or any additional complex operations. Free from the constraints of mention words, VP-MEL supports a broader range of applications. It enables effective entity linking even in situations where users are unfamiliar with the textual language or place greater emphasis on image-based information. To support this new task, we construct a VPWiki dataset based on the existing MEL public datasets, where visual prompts are annotated for each mention within the associated images.
@@
    We construct a high-quality annotated dataset, VPWiki, to enhance the benchmark evaluation of the task. An automated annotation pipeline is proposed to enhance the generation efficiency of the VPWiki dataset.
@@ 2.1 Multimodal Entity Linking
Given the widespread use of image-text content in social media, the integration of both modalities for entity linking is essential. For example, Moon et al. (2018) pioneer the use of images to aid entity linking. Building on this, Adjali et al. (2020) and Gan et al. (2021) construct MEL datasets from Twitter and long movie reviews. Expanding the scope of MEL datasets, Wang et al. (2022c) present a high-quality MEL dataset from Wikinews, featuring diversified contextual topics and entity types. To achieve better performance on these datasets, a multitude of outstanding works in the MEL field (Wang et al., 2022a; Yang et al., 2023; Luo et al., 2023; Shi et al., 2024; Song et al., 2024) emerge, focusing on extracting and interacting with multimodal information. Although multimodal information can enhance entity linking performance, in these methods, text consistently dominates over images.
@@ 2.2 Vision Prompt
Region-specific comprehension in complex visual scenes has become a key research topic in the field of Multimodal Computer Vision. Existing methods typically utilize textual coordinate representations (Zhu et al., 2024; Zhao et al., 2023), learned positional embeddings (Peng et al., 2024; Zhang et al., 2023b; Zhou et al., 2023), or Region of Interest (ROI) features (Zhang et al., 2023b) to anchor language to specific image regions. More recently, Cai et al. (2024) propose a coarse-grained visual prompting solution that directly overlays visual prompts onto the image canvas. In contrast, our VP-MEL provides a fine-grained entity linking method based on visual prompts to reduce reliance on text.
@@ 3 Dataset
As there is no existing MEL dataset that incorporates visual prompts, annotating a high-quality dataset is crucial to enhance the benchmarking of the VP-MEL task.
@@
Annotation Design.
Given an image-text pair and the corresponding mention words, annotators need to: 1) annotate relevant visual prompts in the image based on the mention words; 2) remove samples where the image and mention words are unrelated; 3) re-annotate samples with inaccurate pipeline annotations; 4) annotate the entity type for each example (i.e., Person, Organization, Location, Country, Event, Works, Misc).
Annotation Procedure.
To improve data annotation efficiency, we develop a pipeline that automatically annotates visual prompts in images based on mention words inspired by Li et al. (2024). In the pipeline, the Visual Entailment Module is employed to evaluate and filter out the highly relevant data. Subsequently, the Visual Grounding Module annotates the visual prompts in the images. The details of modules in the pipeline are provided in the Appendix A.6. The annotation team consists of 10 annotators and 2 experienced experts. All annotators have linguistic knowledge and are instructed with detailed annotation principles. Fleiss Kappa score (Fleiss, 1971) of annotators is 0.83, indicating strong agreement among them. We use the Intersection over Union (IoU) metric to assess annotation quality and discard samples with an IoU score below 0.5.
Figure 2: An example from VPWiki and WikiDiverse. The red box in the left image represents the visual prompt annotated for the VP-MEL task. The red text in the right image shows the annotated mention words.
@@
Figure 3: More statistics of VPWiki. (a) Distribution of entity types. (b) Distribution of the number of candidate entities per mention.
@@
Figure 2 illustrates an example from the VPWiki dataset and highlights its differences from traditional MEL data. The VPWiki dataset comprises a total of 12,720 samples, which are randomly split into training, validation, and test sets with an 8:1:1 ratio. Detailed statistics of the VPWiki dataset are provided in Table 1. Additionally, Figure 3 presents the distribution of entity types and the number of candidate entities per mention in the dataset. In Figure 3(a), abbreviations are used to represent each entity type. Meanwhile, Figure 3(b) shows that as the number of candidate entities per mention increases, the task becomes increasingly challenging.
4 Task Formulation
The multimodal knowledge base is constructed by a set of entities =$\mathcal{E}=\{E_{i}\}_{i=1}^{N}$, where each entity is denoted as $E_{i}=(e_{n_{i}},e_{v_{i}},e_{d_{i}},e_{a_{i}})$. Here, $e_{n_{i}}$ denotes the entity name, $e_{v_{i}}$ represents the entity images, $e_{d_{i}}$ corresponds to the entity description, and $e_{a_{i}}$ captures the entity attributes. A mention is denoted as $M_{j}=(m_{s_{j}},m_{v_{j}})$, where $m_{s_{j}}$ represents the sentence and $m_{v_{j}}$ denotes the corresponding image. The related entity of mention $M_{j}$ in the knowledge base is $E_{i}$. The objective of the VP-MEL task is to retrieve the ground truth entity $E_{i}$ from the entity set $\mathcal{E}$ in the knowledge base, based on $M_{j}$.
@@ 5 Methodology
In this section, we describe the proposed FBMEL framework for the VP-MEL task. As illustrated in Figure 4, FBMEL utilizes visual encoder to extract both deep semantic features and shallow texture features, which are enhanced by visual prompts (§5.1). To avoid excessive reliance on visual features, the Detective-VLM module is designed to generate supplementary textual information guided by visual prompts(§5.2), which is then combined with the original text and processed by the text encoder (§5.3). Finally, a similarity score is computed after integrating the visual and textual features(§5.4).
Figure 4: The overall architecture of Feature Balanced Multimodal Entity Linking (FBMEL) framework. The image-text pairs of the Mention and Entity are used together as input. Specifically, Mention Text is the sentence corresponding to Mention Image, while Entity Text consists of Entity Name and Entity Attribute corresponding to the Entity Image in the Knowledge Base.
@@ 5.1 Visual Encoder
To extract image features, we choose pre-trained CLIP model(Dosovitskiy et al., 2021) as our visual encoder. The image $m_{v_{j}}$ of $M_{j}$ is reshaped into n 2D patches. After this, image patches are processed through visual encoder to extract features. The hidden states extracted from $m_{v_{j}}$ by the CLIP visual encoder are represented as $V_{M_{j}}^{l}=\left[v_{[CLS]}^{0};v_{M_{j}}^{1};v_{M_{j}}^{2};...;v_{M_{j}}^{n}\right]\in\mathbb{R}^{(n+1)\times d_{c}}$, where $d_{c}$ denotes the dimension of the hidden state and $l$ denotes the number of layers in the encoder.
@@ 5.2 Detective-VLM
Background: $\{Image\}$ Text: $\{Sentence\}$ Question: In the red box of the image, tell me briefly what is the $\{Entity\ \allowbreak Type\}$, and which $\{Entity\ \allowbreak Name\}$ in the text corresponds to the $\{Entity\ \allowbreak Type\}$? Answer: $\{Entity\ \allowbreak Name\}$$\{Entity\ \allowbreak Type\}$
We utilize VPWiki dataset to design the fine-tuning dataset, where $\{Image\}$ and $\{Sentence\}$ correspond to $m_{v_{j}}$ and $m_{s_{j}}$ in $M_{j}$, respectively. During the inference process, $\{Entity\ \allowbreak Name\}$ and $\{Entity\ \allowbreak Type\}$ need to be generated by VLM. Details of the dataset and Detective-VLM can be found in Appendix A.4.
@@ 6.1 Experimental Settings
All the training and testing are conducted on a device equipped with 4 Intel(R) Xeon(R) Platinum 8380 CPUs and 8 NVIDIA A800-SXM4-80GB GPUs. Detailed experimental settings are provided in Appendix A.1. To comprehensively evaluate the effectiveness of our approach, we compare FBMEL with various competitive MEL baselines and VLM baselines. A detailed introduction of these baselines is provided in the Appendix A.2.
For the VP-MEL task experiments, all approaches are evaluated on the VPWiki dataset. And for the MEL task experiments, all approaches are evaluated on the WikiDiverse (Wang et al., 2022c) dataset. Additional experiments and detailed explanations are provided in the Appendix A.8.
@@ 6.2 Main Results
As shown in Table 2(a), FBMEL significantly outperforms all other methods on VP-MEL task. First, among the VLM methods, LLaVA-1.5 has the smallest performance gap compared to our method, with differences of 5.14%, 1.97%, and 10.58% from FBMEL across the three metrics, respectively. Even so, considering the substantial data and computational resources required to train LLaVA, FBMEL still demonstrates an obvious advantage. Second, there is a notable performance gap between MEL methods and FBMEL. MEL methods struggle with effective entity linking in scenarios where mention words are absent, underscoring their limitations and the robustness of our approach.
@@
Table 2(b) presents the experimental results comparing FBMEL with other methods on MEL dataset. During the testing process, the Detective-VLM is removed from FBMEL, and its output is replaced with mention words from WikiDiverse. The performance of FBMEL is slightly behind MIMIC and MELOV, with differences of only 0.27%, 0.28%, and 0.07% across the three metrics. This is expected, as FBMEL places greater emphasis on image features while reducing reliance on text features. Furthermore, the absence of visual prompts in the images limits the ability of FBMEL to fully utilize visual information.